OpenStudy (anonymous):

2/3=t/6 -what is the value of t 2/3=5/r -what is the value r

OpenStudy (anonymous):

2/3=t/6 -what is the value of t t= (2/3) *6= 4 2/3=5/r -what is the value r r= 5/(2/3) =15/2